OpenStudy (anonymous):

what is the integral of (lnx)^2 dx

OpenStudy (anonymous):

Try substituting u=ln(x).

OpenStudy (anonymous):

Let u = [ ln(x) ]^2. dv = dx. du = 2ln(x) (1/x) dx. v = x Recall the formula of integration by parts: uv - Integral ( v du ) x [ln(x)]^2 - Integral ( x (2 ln(x)) (1/x) dx ) And we get some cancellations in the integral, particular with x and (1/x). x [ln(x)]^2 - Integral ( 2 ln(x) dx ) And let's factor out the 2 from the integral. x [ln(x)]^2 - 2 * Integral ( ln(x) dx ) From here, you would use integration by parts again. Let u = ln(x). dv = dx. du = (1/x) dx. v = x x [ln(x)]^2 - 2[ x ln(x) - Integral ( x(1/x) dx ) Distribute the -2, to get x [ln(x)]^2 - 2x ln(x) + 2 * Integral ( x(1/x) dx ) Simplify the integral, x [ln(x)]^2 - 2x ln(x) + 2 * Integral ( 1 dx ) And now it's an easy integral. With no more integrals to solve, don't forget to add a constant. x [ln(x)]^2 - 2x ln(x) + 2x + C Factor the common monomial x from the three terms. x ( [ ln(x)]^2 - 2 ln(x) + 2 ) + C